OpenStudy (anonymous):

3y + 8 = =5 = 10y -3y

OpenStudy (goten77):

want to simplify? or solve for y? or wahts going on

OpenStudy (anonymous):

i need to solve the equation an i dont know how too

OpenStudy (anonymous):

...

OpenStudy (anonymous):

wats the .... for

OpenStudy (goten77):

ugh what you have is 3y+8 = 5 = 10y-3y is this correct?

OpenStudy (anonymous):

yea but its have to be a negative 5

OpenStudy (goten77):

oh thats what u meant

OpenStudy (anonymous):

yes

OpenStudy (goten77):

so your saying 3y+8= -5 = 10y-3y

OpenStudy (anonymous):

- 10 but yes thats what i mean

OpenStudy (goten77):

um how about you rerwite exactly what you want... im not sure now

OpenStudy (goten77):

or exactly what it shows

OpenStudy (goten77):

ima assume 3y+8 = -5 -10y -3y ???????????